Firm A Revenues and Costs Thousands of Dollars Firm B Revenues and Costs Thousands of Dollars 280 Total otal Costs Total Costs 28 Total 240 20 - 160 120 240 200 160 120 80 40 Breakeven Point (30 240) Breakeven Point 122222, 188 889) l Fixed Costs Foed Costs 80 40 0 10 20 30 40 50 60 Unts (Thousands) 0 10 20 30 40 50 60 Units (Thousands) a. Given the graphs above, calculate the total foxed costs, variable costs per unit, and sales price for Firm A. Firm Bs fixed costs are $120,000, its variable costs per unit are $4, and its sales price is 8 per unit. Round your answers to two decimal places Fixed costs are s Variable costs per unit are s 4.90 Sales price per unit is $ 8.50 b. Which firm has the higher operating leverage at any given level of sales? c. At what sales level, in units, do both firms eam the same operating profit? Round intermediate calculations to 2 decimal places. Round your answer to the nearest whole number 4444.44 units.
Answer #1

A) As from the graph for firm A it is shown that fxed cost line is starting from $80000 and also at this point on y axis, the cost line starts at $80,000.

variable cost

Breakeven point for firm A is (22.22, 188.89) means that total cost is 188.89 for 22.22

total cost = fixed cost + variable cost per unit* no of unit

188.89 = 80 +22.22* x

108.89 = 22.22x

x = 4.9

Sales per unit

at breakeven,

Sales = total fixed cost + total variable cost=

Sales = (80 +4.9*22.22)

sales = 188.89

Sales per unit = 188.89/22.22 = 8.5

B) operating leverage is define as the change in operating profit/sales. As from graph it is shown that for firm B, cost and revenue line are more diverting and steeper tha firm A. More diverting means more far apart. So, firm B is corrct option

C) When both the firm has same opeating profit,

firm A operating profit = firm B operating profit

firm A(Sales - fixed cost - variable cost) = firm B(Sales - fixed cost - variable cost)

8.5x - 80 - 4.9x = 8x-120-4x

40 = 0.9x -0.5x = 0.4x

x = 100 thousand units